### v1

@@ 1 Introduction
PoseNet [13] is the first study on the deep learning-based end-to-end localization method. It estimates the six-dimensional pose directly from sensor data (image) using neural networks in an end-to-end manner. Since PoseNet, the following studies have been conducted based on image [25], 3D point cloud [28, 30, 17, 16], inertial information [9], and the fusion of image and 2D point cloud [15]. These end-to-end localization methods are known to be more robust against sensor data variations such as noise and illumination. However, it is difficult to use them solely because they generally provide relatively higher localization errors compared to the matching-based localization methods. [11] utilized PoseNet to get an initial pose for a particle filter-based localization. However, they overlooked how much confidence we can have in the output of PoseNet. If the PoseNet output with a high localization error is used as an initial pose, the navigation system may not be operated.
In this paper, we introduce an uncertainty-aware rejection mechanism to improve localization accuracy. In recent, [16] proposed measuring uncertainty from the outputs of their proposed method DiffLoc, which is a diffusion model for localization to estimate the 6D pose of a 3D LiDAR sensor from a point cloud. Their experiments showed that the measured uncertainty is highly correlated with localization error. We go one step further beyond [16]. We leverage uncertainty as a threshold. It allows the rejection of localization results with high errors, and we can trust the non-rejected localization results through the rejection strategy. Experiments using our datasets collected by a 2D LiDAR and a camera demonstrate that the uncertainty-based rejection effectively reduces position and orientation errors. More specifically, our uncertainty-based rejection method reduces the position error by up to 69.4% and the orientation error by 73.3%. In particular, experimental results show that our strategy can exclude the outputs with significant position and orientation errors. This means that the passed outputs are reliable enough for the results of global localization. To the best of our knowledge, this is the first study that systematically utilizes uncertainty-based thresholds to reject unreliable localization results predicted in an end-to-end manner and improves the accuracy and confidence of pose estimates.
@@ 2.1 End-to-end localization
Visual localization estimates the current pose using only images captured by a camera in indoor or outdoor environments. Initially, convolutional neural networks (CNNs) are primarily leveraged to extract salient features. However, recent studies have introduced various techniques in their models. [13] developed a CNN-based 6-DoF pose regression model that allows localization without relying on feature matching or keyframes. Meanwhile, [25] employed the self-attention technique [24] improve the accuracy of the end-to-end localization. Moreover, Transformer architectures [24] have been utilized for end-to-end camera localization as well [18, 22]. In recent, [27] integrated CNNs, self-attention, and long short-term memory (LSTM) modules in a unified architecture to extract static features, which can lead to more effective 6-DoF pose estimation compared to using dynamic features.
LiDAR-based end-to-end localization leverages 3D structural information, making it robust in textureless environments. [28] introduced the first LiDAR-based 6-DoF pose regression model, enhancing feature learning with self-attention mechanism. [30] proposed a deep neural network for pose regression that consists of two modules: a universal encoder for scene feature extraction and a regressor for pose estimation. They also demonstrated the relationship between the regression capability and the number of hidden units in the regression module. [29] introduced additional classification headers alongside the original regression headers, together with a feature aggregation module based on temporal attention for spatial and temporal constraints. [10] presented a self-supervised learning approach utilizing a Transformer-based backbone for LiDAR-based end-to-end localization. Also, SGLoc [17] enhanced pose regression accuracy by incorporating scene geometry encoding. Lastly, [16] enhanced accuracy further by applying an iterative denoising process based on a diffusion model to the pose regression.
Some studies have combined complementary information from multiple modalities to improve the robustness of localization. For instance, [14] proposed leveraging both visual and LiDAR features to achieve more accurate and robust place recognition. [26] developed a vision-assisted LiDAR localization method that effectively utilizes visual information to address issues related to 2D LiDAR-based localization drift. Additionally, [20] incorporated a fisheye camera together with a 2D LiDAR system to enhance localization fault detection. However, the methods mentioned above do not fall under the category of end-to-end localization techniques. FusionLoc [15] is an end-to-end localization method that utilizes multi-modality. In this study, we present a FusionLoc-based approach to make localization more reliable by rejecting network outputs with significant errors.
@@ 2.2 Uncertainty quantification
Uncertainty quantification is a well-established topic in pattern recognition and machine learning. While it did not receive much attention during the early stages of the deep learning revolution—especially in comparison to efforts to enhance the accuracy of deep learning algorithms—its importance is becoming increasingly recognized, particularly in safety-critical applications. A Bayesian approach is one of the most comprehensive frameworks for managing uncertainty. However, developing and implementing a Bayesian deep neural network for regression tasks is very challenging because it is often impractical to determine posterior probabilities accurately. Fortunately, [7] presented dropout as an alternative to Bayesian approximation. Additionally, [12] proposed using Monte Carlo (MC) dropout to quantify both aleatoric and epistemic uncertainties in regression tasks, such as pixel-wise depth estimation.
After the groundbreaking studies, the following researchers focused on network calibration, which aims to align estimated uncertainty values with empirical results. Kuleshov et al. introduced a simple, algorithm-agnostic method inspired by Platt scaling [21]. [6] utilized the maximum mean discrepancy, viewing calibration as a form of distribution matching. Similarly, [2] applied the f-divergence with the same perspective. In another approach, [31] proposed an auxiliary network branch to estimate uncertainty alongside the main branch used for the original regression task. This method is similar to the work of [5], which also employs additional network branches to estimate uncertainty or confidence in classification problems. For more details on uncertainty qualification in deep neural networks, refer to [1, 8].
However, none of the studies mentioned addressed the localization problem. [4] recently proposed a method for quantifying uncertainty in visual localization. However, their approach estimates the pose of a query image through keypoint matching instead of an end-to-end method. In contrast, [16] suggested an end-to-end localization approach using a diffusion model. However, their work primarily focused on the relationship between quantified uncertainties (variance) and positional errors, lacking qualitative experimental results. Unlike [4] and [16], this study aims to quantify uncertainty in the results of an end-to-end localization method. We will also demonstrate how this quantification can effectively reject network outputs with significant errors, ultimately improving the localization performance.
@@ 3.1 FusionLoc
FusionLoc [15] is a deep learning-based robot localization method that combines RGB images and 2D range data to improve the localization accuracy by leveraging the strengths of both sensors. Specifically, FusionLoc predicts the robot’s 3-DoF pose, including planar position and orientation, using an image $\mathbf{I}$ and 2D range data $\mathbf{S}$ as inputs as the following:
@@
where $\hat{\mathbf{p}}=[\hat{x},\hat{y}]$ is the 2D coordinates of the robot position, and $\hat{\mathbf{q}}=[\cos\hat{\theta},\sin\hat{\theta}]$ corresponds to the robot orientation. The method computes an image feature from the input image using a feature extractor from AtLoc [25], while it calculates point features from the input range data using a different feature extractor from PointLoc [28]. To enhance the interaction between these two modalities, multi-head self-attention [24] is employed. This approach enables more effective multi-modality fusion than traditional methods such as concatenation or addition of the image and point features. Lastly, the output of the multi-head self-attention block is passed through the regression block. The regression block has two branches responsible for position and orientation. Each branch consists of successive MLPs.
@@ 3.2 Measureing uncertainty
Uncertainty can be divided into two categories: epistemic uncertainty and aleatoric uncertainty [12]. Epistemic uncertainty arises from limitations in the model’s knowledge or training process, typically due to insufficient data. This uncertainty can be reduced by incorporating additional training data or enhancing the model architecture. In contrast, aleatoric uncertainty stems from sensor noise, measurement errors, or inherent randomness in the data collection procedure. Unlike epistemic uncertainty, aleatoric uncertainty cannot be eliminated through additional training, as it originates from data sensing. Both types of uncertainty can be estimated using Bayesian neural networks (BNNs). BNNs treat model weights as probabilistic distributions to quantify uncertainty. However, computing the exact posterior distribution in high-dimensional spaces is almost impractical. In this study, we use MC dropout to approximate Bayesian inference and provide uncertainty estimation.
Let us consider a regression task with $N$ data pairs of input $\mathbf{x}$ and output $y$, i.e., $\left(\mathbf{x}_{i},y_{i}\right)_{i=1}^{N}$. To quantify aleatoric and epistemic uncertainties in this task, we consider a BNN model $\mathbf{f}$ to infer the posterior distribution. From an input $\mathbf{x}$, it provides a model output $\hat{y}$ together with a variance $\hat{\sigma}^{2}$ of the aleatoric uncertainty. In contrast, to estimate the epistemic uncertainty, we employ the MC dropout to approximate the posterior over the model. By representing the model weights as $\hat{\mathbf{W}}$ from the approximate posterior, the model provides both the predictive mean and the variance, i.e., $[\hat{y},\hat{\sigma}^{2}]=\mathbf{f}^{\hat{\mathbf{W}}}(\mathbf{x})$. The objective function of learning the model can be defined without the regularization term as the following:
@@ 3.3 Uncertainty-aware localization
As mentioned above, we replace $\hat{\sigma}_{\mathbf{p}}^{2}$ and $\hat{\sigma}_{\mathbf{q}}^{2}$ by $s_{\mathbf{p}}=\log\hat{\sigma}_{\mathbf{p}}^{2}$ and $s_{\mathbf{q}}=\log\hat{\sigma}_{\mathbf{q}}^{2}$ for computational stability. Also, we measure the epistemic uncertainty by applying the MC dropout to the output of the self-attention block mentioned above. Given $N$ training samples $\{(\mathbf{I}_{i},\mathbf{S}_{i},\mathbf{p}_{i},\mathbf{q}_{i})\}_{i=1}^{N}$, the loss function can be defined as the following:
@@ 4.1 Datasets
Position (m) Min 0.015 0.018 0.002 0.016 0.005 0.005 0.007 0.008 0.007 0.007 0.003 0.005
Median 0.104 0.099 0.097 0.095 0.056 0.057 0.057 0.057 0.114 0.105 0.100 0.094
Max 0.967 0.557 0.503 0.458 0.339 0.277 0.249 0.217 3.973 1.168 0.618 0.605
Mean 0.136 0.119 0.119 0.110 0.064 0.063 0.063 0.062 0.214 0.143 0.135 0.128
Orientation () Min 0.013 0.13 0.010 0.013 0.012 0.013 0.009 0.008 0.004 0.004 0.005 0.013
Median 2.693 2.434 2.469 2.718 1.127 1.018 0.942 0.857 2.423 2.160 2.083 1.938
Max 73.406 20.407 14.891 13.040 25.895 14.721 9.914 7.322 168.234 48.723 16.330 11.795
Mean 4.875 3.360 2.933 2.716 1.989 1.621 1.428 1.299 6.177 2.960 2.507 2.290
@@ 4.2 Evaluation
Table 2 illustrates the changes in position and orientation errors under varying uncertainty thresholds applied to the TheGardenParty, ETRI, SusungHotel datasets. The results demonstrate consistent reductions in mean position and orientation errors across all rejection thresholds (90%, 80%, and 70%). In the TheGardenParty dataset, applying a 70% uncertainty threshold led to a reduction in the mean position error by as much as 19.1% and a decrease in the mean orientation error by up to 44.3%. For the ETRI dataset under the same conditions, the reductions were 3.1% for position error and 34.7% for orientation error. Notably, the effectiveness of the uncertainty-based rejection strategy was also observed in the SusungHotel datasets. For the SusungHotel dataset, the mean position error decreased from 0.214 m to 0.128 m, while the mean orientation error was reduced from 6.177 to 2.290, reflecting improvements of 40.2% and 62.9%, respectively. These findings indicate that the TheGardenParty and SusungHotel datasets have more outliers and noise, which leads to more significant performance improvements with the uncertainty-based rejection strategy. In contrast, the ETRI dataset, likely collected in a more stable environment, shows lower errors even without applying the rejection method, resulting in relatively less impact from this approach. Significantly, in the TheGardenParty dataset, the maximum position error decreased from 0.967 m to 0.458 m, and the maximum orientation error significantly dropped from 73.406 to 13.040. This demonstrates the effectiveness of the rejection method in addressing extreme localization errors. These results indicate our end-to-end localization with the uncertainty-based rejection method can be utilized as a global localization solution, which can also provide a reliable initial pose to conventional localization modules, e.g., the adaptive Monte Carlo localization, which is known as AMCL.

### v2

@@ 1 Introduction
PoseNet [13] is the first study on the deep learning-based end-to-end localization method. It estimates the six-dimensional pose directly from sensor data (image) using neural networks in an end-to-end manner. Since PoseNet, the following studies have been conducted based on image [26], 3D point cloud [29, 30, 17, 18], inertial information [9], and the fusion of image and 2D point cloud [16]. These end-to-end localization methods are known to be more robust against sensor data variations such as noise and illumination. However, it is difficult to use them solely because they generally provide relatively higher localization errors compared to the matching-based localization methods. Jo and Kim [11] utilized PoseNet to get an initial pose for a particle filter-based localization. However, they overlooked how much confidence we can have in the output of PoseNet. If the PoseNet output with a high localization error is used as an initial pose, the navigation system may not be operated.
In this paper, we introduce an uncertainty-aware rejection mechanism to improve localization accuracy. In recent, Li et al. [18] proposed measuring uncertainty from the outputs of their proposed method DiffLoc, which is a diffusion model for localization to estimate the 6D pose of a 3D LiDAR sensor from a point cloud. Their experiments showed that the measured uncertainty is highly correlated with localization error. We go one step further beyond [18]. We leverage uncertainty as a threshold. It allows the rejection of localization results with high errors, and we can trust the non-rejected localization results through the rejection strategy. Experiments using our datasets collected by a 2D LiDAR and a camera demonstrate that the uncertainty-based rejection effectively reduces position and orientation errors. More specifically, our uncertainty-based rejection method reduces the position error by up to 69.4% and the orientation error by 73.3%. In particular, experimental results show that our strategy can exclude the outputs with significant position and orientation errors. This means that the passed outputs are reliable enough for the results of global localization. To the best of our knowledge, this is the first study that systematically utilizes uncertainty-based thresholds to reject unreliable localization results predicted in an end-to-end manner and improves the accuracy and confidence of pose estimates.
@@ 2.1 End-to-end localization
Visual localization estimates the current pose using only images captured by a camera in indoor or outdoor environments. Initially, convolutional neural networks (CNNs) are primarily leveraged to extract salient features. However, recent studies have introduced various techniques in their models. Kendall et al. [13] developed a CNN-based 6-DoF pose regression model that allows localization without relying on feature matching or keyframes. Meanwhile, Wang et al. [26] employed the self-attention technique [25] improve the accuracy of the end-to-end localization. Moreover, Transformer architectures [25] have been utilized for end-to-end camera localization as well [19, 23]. In recent, Wang et al. [28] integrated CNNs, self-attention, and long short-term memory (LSTM) modules in a unified architecture to extract static features, which can lead to more effective 6-DoF pose estimation compared to using dynamic features.
LiDAR-based end-to-end localization leverages 3D structural information, making it robust in textureless environments. Wang et al. [29] introduced the first LiDAR-based 6-DoF pose regression model, enhancing feature learning with self-attention mechanism. Yu et al. [30] proposed a deep neural network for pose regression that consists of two modules: a universal encoder for scene feature extraction and a regressor for pose estimation. They also demonstrated the relationship between the regression capability and the number of hidden units in the regression module. Yu et al. [31] introduced additional classification headers alongside the original regression headers, together with a feature aggregation module based on temporal attention for spatial and temporal constraints. Ibrahim et al. [10] presented a self-supervised learning approach utilizing a Transformer-based backbone for LiDAR-based end-to-end localization. Also, SGLoc [17] enhanced pose regression accuracy by incorporating scene geometry encoding. Lastly, Li et al. [18] improved accuracy further by applying an iterative denoising process based on a diffusion model to the pose regression.
Some studies have combined complementary information from multiple modalities to improve the robustness of localization. For instance, Lai et al. [15] proposed leveraging both visual and LiDAR features to achieve more accurate and robust place recognition. Wang et al. [27] developed a vision-assisted LiDAR localization method that effectively utilizes visual information to address issues related to 2D LiDAR-based localization drift. Additionally, Nakamura et al. [21] incorporated a fisheye camera together with a 2D LiDAR system to enhance localization fault detection. However, the methods mentioned above do not fall under the category of end-to-end localization techniques. FusionLoc [16] is an end-to-end localization method that utilizes multi-modality. In this study, we present a FusionLoc-based approach to make localization more reliable by rejecting network outputs with significant errors.
@@ 2.2 Uncertainty quantification
Uncertainty quantification is a well-established topic in pattern recognition and machine learning. While it did not receive much attention during the early stages of the deep learning revolution—especially in comparison to efforts to enhance the accuracy of deep learning algorithms—its importance is becoming increasingly recognized, particularly in safety-critical applications. A Bayesian approach is one of the most comprehensive frameworks for managing uncertainty. However, developing and implementing a Bayesian deep neural network for regression tasks is very challenging because it is often impractical to determine posterior probabilities accurately. Fortunately, Gal and Ghahramani [7] presented dropout as an alternative to Bayesian approximation. Additionally, Kendall and Gal [12] proposed using Monte Carlo (MC) dropout to quantify both aleatoric and epistemic uncertainties in regression tasks, such as pixel-wise depth estimation.
After the groundbreaking studies, the following researchers focused on network calibration, which aims to align estimated uncertainty values with empirical results. Kuleshov et al. [14] introduced a simple, algorithm-agnostic method inspired by Platt scaling [22]. Cui et al. [6] utilized the maximum mean discrepancy, viewing calibration as a form of distribution matching. Similarly, Bhatt et al. [2] applied the $f$-divergence with the same perspective. In another approach, Yu et al. [32] proposed an auxiliary network branch to estimate uncertainty alongside the main branch used for the original regression task. This method is similar to the work of [5], which also employs additional network branches to estimate uncertainty or confidence in classification problems. For more details on uncertainty qualification in deep neural networks, refer to [1, 8].
However, none of the studies mentioned above addressed the localization problem. Chen et al. [4] recently proposed a method for quantifying uncertainty in visual localization. However, their approach estimates the pose of a query image through keypoint matching rather than an end-to-end method. In contrast, Li et al. [18] suggested an end-to-end localization approach using a diffusion model. However, their work primarily focused on the relationship between quantified uncertainties (variance) and positional errors, lacking qualitative experimental results. Unlike [4] and [18], this study aims to quantify uncertainty in the results of an end-to-end localization method. We will also demonstrate how this quantification can effectively reject network outputs with significant errors, ultimately improving the localization performance.
@@ 3.1 FusionLoc
FusionLoc [16] is a deep learning-based robot localization method that combines an RGB image and 2D range data to improve the localization accuracy by leveraging the strengths of both sensors. Specifically, FusionLoc predicts the robot’s 3-DoF pose, including planar position and orientation, using an image $\mathbf{I}$ and 2D range data $\mathbf{S}$ as inputs as the following:
@@
where $\hat{\mathbf{p}}=[\hat{x},\hat{y}]$ is the 2D coordinates of the robot position, and $\hat{\mathbf{q}}=[\cos\hat{\theta},\sin\hat{\theta}]$ corresponds to the robot orientation. The method computes an image feature from the input image using a feature extractor from AtLoc [26], while it calculates point features from the input range data using a different feature extractor from PointLoc [29]. To enhance the interaction between these two modalities, multi-head self-attention [25] is employed. This approach enables more effective multi-modality fusion than traditional methods such as concatenation or addition of the image and point features. Lastly, the output of the multi-head self-attention block is passed through the regression block. The regression block has two branches responsible for position and orientation. Each branch consists of successive MLPs to provide the position $\hat{\mathbf{p}}$ and the orientation $\hat{\mathbf{q}}$.
@@ 3.2 Measureing uncertainty
Uncertainty can be divided into two categories: epistemic uncertainty and aleatoric uncertainty [12]. Epistemic uncertainty arises from limitations in the model’s knowledge or training process, typically due to insufficient data. This uncertainty can be reduced by incorporating additional training data or enhancing the model architecture. In contrast, aleatoric uncertainty stems from sensor noise, measurement errors, or inherent randomness in the data collection procedure. Unlike epistemic uncertainty, aleatoric uncertainty cannot be eliminated through additional training, as it originates from data sensing. Both types of uncertainty can be estimated using Bayesian neural networks (BNNs). BNNs treat model weights as probabilistic distributions to quantify uncertainty. However, computing the exact posterior distribution in high-dimensional spaces is almost impractical. In this study, we use MC dropout [12] to approximate Bayesian inference and provide uncertainty estimation.
Let us consider a regression task with $N$ data pairs of input $\mathbf{x}$ and output $y$, i.e., $\{\left(\mathbf{x}_{i},y_{i}\right)\}_{i=1}^{N}$. To quantify aleatoric and epistemic uncertainties in this task, we consider a BNN model $\mathbf{f}$ to infer the posterior distribution. From an input $\mathbf{x}$, it provides a model output $\hat{y}$ together with a variance $\hat{\sigma}^{2}$ of the aleatoric uncertainty. In contrast, to estimate the epistemic uncertainty, we employ the MC dropout to approximate the posterior over the model. By representing the model weights as $\hat{\mathbf{W}}$ from the approximate posterior, the model provides both the predictive mean and the variance, i.e., $[\hat{y},\hat{\sigma}^{2}]=\mathbf{f}^{\hat{\mathbf{W}}}(\mathbf{x})$. The objective function of learning the model can be defined without the regularization term as the following:
@@ 3.3 Uncertainty-aware localization
Fig. 1 shows the pipeline of our model. As mentioned above, we replace $\hat{\sigma}_{\mathbf{p}}^{2}$ and $\hat{\sigma}_{\mathbf{q}}^{2}$ by $s_{\mathbf{p}}=\log\hat{\sigma}_{\mathbf{p}}^{2}$ and $s_{\mathbf{q}}=\log\hat{\sigma}_{\mathbf{q}}^{2}$ for computational stability. Also, we measure the epistemic uncertainty by applying the MC dropout to the output of the self-attention block mentioned above. Given $N$ training samples $\{(\mathbf{I}_{i},\mathbf{S}_{i},\mathbf{p}_{i},\mathbf{q}_{i})\}_{i=1}^{N}$, the loss function can be defined as the following:
@@ 4.1 Datasets
Position (m) Min 0.072 0.074 0.069 0.098 0.023 0.025 0.026 0.031 0.022 0.028 0.031 0.019
Median 0.573 0.567 0.552 0.523 0.289 0.298 0.300 0.298 0.659 0.561 0.477 0.424
Max 5.559 5.288 4.317 4.192 1.610 1.613 1.287 1.490 66.392 66.279 59.829 59.753
Mean 0.776 0.738 0.712 0.682 0.324 0.326 0.327 0.327 4.262 2.788 2.211 1.575
Orientation () Min 0.0204 0.017 0.003 0.015 0.013 0.012 0.008 0.010 0.004 0.008 0.009 0.008
Median 2.672 2.544 2.512 2.299 1.127 1.052 0.978 0.902 2.732 2.466 2.168 2.020
Max 72.738 54.508 49.384 48.397 26.224 16.182 15.569 11.024 179.385 174.738 173.323 127.828
Mean 4.880 3.783 3.358 3.126 1.990 1.66086 1.52688 1.409 13.711 9.008 5.861 4.329
@@ 4.2 Evaluation
Table 2 illustrates the changes in position and orientation errors under varying uncertainty thresholds applied to the TheGardenParty, ETRI, SusungHotel datasets. The results demonstrate consistent reductions in mean position and orientation errors across all rejection thresholds (90%, 80%, and 70%). In the TheGardenParty dataset, applying a 70% uncertainty threshold led to a reduction in the mean position error by as much as 12.1% (from 0.776 m to 0.682 m) and a decrease in the mean orientation error by up to 36.0% (from 4.880 to 3.126).. For the ETRI dataset under the same conditions, the reductions were 0.6and 29.2% for orientation error (from 1.990 to 1.409). Notably, the effectiveness of the uncertainty-based rejection strategy was also observed in the SusungHotel datasets. For the SusungHotel dataset, the mean position error decreased from 4.262 m to 1.575 m, while the mean orientation error was reduced from 13.711 to 4.329, reflecting improvements of approximately 63.0These findings indicate that the TheGardenParty and SusungHotel datasets have more outliers and noise, which leads to more significant performance improvements with the uncertainty-based rejection strategy. In contrast, the ETRI dataset, likely collected in a more stable environment, shows lower errors even without applying the rejection method, resulting in relatively less impact from this approach. Significantly, in the TheGardenParty dataset, the maximum position error decreased from 5.559 m to 4.192 m, and the maximum orientation error significantly dropped from 72.738 to 48.397. This demonstrates the effectiveness of the rejection method in addressing extreme localization errors. These results indicate our end-to-end localization with the uncertainty-based rejection method can be utilized as a global localization solution, which can also provide a reliable initial pose to conventional localization modules, e.g., the adaptive Monte Carlo localization, which is known as AMCL.